OpenStudy (anonymous):

can any help me with the deriv. of 16w+6((310/w))+3w???? and the critical ppoints/// i cant get a right answer

OpenStudy (anonymous):

\[16w+6(310/w)+3w\] right?

OpenStudy (anonymous):

yea

OpenStudy (dumbcow):

first add like terms then find derivative of each term using the power rule

OpenStudy (anonymous):

would i multiply 6by 310?

OpenStudy (dumbcow):

yes

OpenStudy (dumbcow):

remember 1/w = w^-1

OpenStudy (anonymous):

so i got the function to be derived... (1860/w)+19w?,,

OpenStudy (dumbcow):

correct

OpenStudy (anonymous):

isnt that the quotient rule then ? i got (1860/w^2)+19

OpenStudy (dumbcow):

you could use the quotient rule, however be careful about the sign it should be -(1860/w^2) +19 (f/g)' = f'g - g'f/g^2

OpenStudy (dumbcow):

now set it equal to zero and solve for w

OpenStudy (dumbcow):

you should get w^2 = 1860/19

OpenStudy (anonymous):

right igot that ...

OpenStudy (anonymous):

+_9.894176916

OpenStudy (dumbcow):

ok good those are your critical points

OpenStudy (anonymous):

ok thanks you lots...